OpenStudy (anonymous):

I have a chemistry question. Convert 100.00degreesC to kelvins Help? : ^ )

OpenStudy (anonymous):

C+273.15=K

OpenStudy (anonymous):

you add 273.15 to your celsius temperature to get the kelvin temperature

OpenStudy (anonymous):

373.15

OpenStudy (anonymous):

he was saying the answer i think

OpenStudy (anonymous):

oh woops you wee answering his question sorry

OpenStudy (anonymous):

who's on first?

OpenStudy (anonymous):

yes 373.15 is the answer

OpenStudy (anonymous):

no. c to K conversion is c+100=k so 273.15c is 373.15K

OpenStudy (anonymous):

yep sorry I thought you were saying the conversions should be +373.15

OpenStudy (anonymous):

i though who's on second?

OpenStudy (anonymous):

i don't know is on second

OpenStudy (anonymous):

Thank you! That really helped!

OpenStudy (anonymous):

well then who's on third?

OpenStudy (anonymous):

What is the process of converting F degrees to C degrees?

OpenStudy (anonymous):

As in 68 degrees F to C degrees

OpenStudy (anonymous):

c = (5/9) (f-32)

OpenStudy (anonymous):

and f=c(5/9)+32

OpenStudy (anonymous):

so you would put 68 in place of F and then ..

OpenStudy (anonymous):

so 68 degrees f is 20 degrees c

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

SO it would l\[68 degreesf=20degrees c(5/9)+32?\]

OpenStudy (anonymous):

yes, then subtract 32, and divide by (9/5)

OpenStudy (anonymous):

why do you subtract 32? I thought you add? And isnt it 5/9? sorry.

OpenStudy (anonymous):

if converting 68F to C?

OpenStudy (anonymous):

So to convert F to C the setup is F=c(5/9) + 32?

OpenStudy (anonymous):

here, ill write them both for you. C=(F-32) x (5/9) F=C x (9/5)+32 if you know the 68 degrees F then you would set it up like this: 68=C x (9/5)+32

OpenStudy (anonymous):

and vice versa for C to F conversion

OpenStudy (anonymous):

Ok thank you.: ^ ) For C its minus 32?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

no worries